OpenStudy (anonymous):

How many solutions does the equation have? 2(2x+5)=4(x+3) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

/@carsonce

OpenStudy (radar):

If you can't solve it, then I would say no solution.

OpenStudy (texaschic101):

2(2x + 5) = 4(x+3) --- distribute through the parenthesis 4x + 10 = 4x + 12 -- subtract 4x from both sides 10 = 12.....incorrect....this means no solution if it would have ended like : 2 = 2 or something equal to each other...then it would be infinite solutions. if it would have ended like : x = 4.....then that means only 1 solution

OpenStudy (texaschic101):

questions ?

OpenStudy (anonymous):

No, thank you :)